OpenStudy (anonymous):

In the early months of some year one site added 0.5 million new accounts everyday.at this rate,how many days would be needed to add 30 million new accounts

OpenStudy (anonymous):

30 million = x * 0.5 million

OpenStudy (anonymous):

So do I multiple 30 million and 0.5

OpenStudy (anonymous):

wait can u give me a min to figure it out then ill explain it to u :)

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

sorry its taking me forever but i did this long time ago i am trying to remember

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):



OpenStudy (anonymous):

30=1+0.4

OpenStudy (anonymous):

@uri @TheSmartOne

OpenStudy (anonymous):

sorry i mean 0.5

OpenStudy (anonymous):

what's the name of the this topic

OpenStudy (anonymous):

Ok so do multiply or divde

OpenStudy (anonymous):

It's linear equations

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

0.5 mil acounts in 1 day how many in 2 days? 2*0.5 right? how many in 3 days? 3*0.5 right? how many in x days? x*0.5 right? Now think of it this way. x is days that it will take to add 30 milion. We can write it in numbers like this x*0.5=30 so, just solve for x.

OpenStudy (anonymous):

Thank you